OpenStudy (anonymous):

how do you graph the system of linear inequalities 2x-2y<6 x-y<9

OpenStudy (amistre64):

ill draw you a picture, but in essense, an inequality is all the points in a given area as opposed to all the points on a given line.

OpenStudy (amistre64):

the lines are y =x-9; and y=x-3 if we do it normally... the point (0,0) is used to calibrate the "area" of the graph; since it is not direclt on a line we can see if it is inside the area or outsid ethe area...

OpenStudy (amistre64):

that whole area that I was doing green will be the "graph" of these inequalities. It important to notice that the lines themsleves are not to be included as a solution because your inequalities are not "equal to"......